# On the Evolution of Virulence in Vector-borne Diseases

### v1

@@ Discussion
The theoretical framework in this study shows that interactions between vector and host transmission position vector-borne diseases within a wide range of virulence states that enable transmission. The vector component may amplify the reproduction number increasing the evolutionary viability for the pathogen and may widen the range of mortality values in which this viability happens. This can lead to increased virulence in hosts, even if it does not correspond necessarily to the highest reproduction numbers. The varying levels of virulence can have serious consequences; for example, yellow fever is known to have a high fatality rate (Douam and Ploss,, 2018). These findings highlight the need for heightened attention to vector-borne diseases as potential emergent threats with pandemic potential.
@@
Geographical and climate conditions may restrict vector distribution, meaning they might not be present globally. de Angeli Dutra et al., (2022) also discuss the possibility of higher virulence in vector-host transmission despite the pathogen needing to overcome different immunological systems and the fitness costs to vectors. However, where vectors are present, there is a significant risk of pathogen emergence or adaptation. Therefore, vector-borne diseases pose a high risk for focused surveillance of infectious diseases.
@@
The model included a growth rate of pathogens within hosts and vectors in which diminishing transmission rate over virulence levels are crucial. This is biologically plausible since the pathogen will grow in hosts and vectors where there is a finite number of potential cells or tissues to infect. Previous theoretical frameworks have also considered diminishing returns (Dieckmann et al.,, 2002). The tradeoff theory has been discussed by several authors including Ewald, (2004), and a model formulation by Bolker et al., (2009). Recent empirical observations also pointed to diminishing rates for virus titers in SARS-CoV-2 (Mautner et al.,, 2022) in patients and DENV-2 in mosquitoes (Salazar et al.,, 2007). Nevertheless, several authors have also examined the challenges of reproducing such effects in various disease contexts, both experimentally and observationally. Alizon et al., (2009) provide a review discussing why it is often challenging to evaluate the tradeoff hypothesis empirically. Doumayrou et al., (2013) found evidence for the tradeoff hypothesis with plant pathogens, and Jensen et al., (2006) found evidence in experiments with a castrating bacterium.
@@
The initial formulation of the model is quite general to provide intuition on the effects of parameters on the basic reproduction number. However, a few assumptions were applied to obtain some of the theoretical results. Typically, the model assumes that the distributions of general mortality and recovery rates follow an exponential pattern. This assumption simplifies the model, making it easier to analyze and draw meaningful conclusions. The formula for the basic reproduction number in this study is arguably more general than the classical MacDonald and Ross formulation (Smith et al.,, 2012), which is typically used in traditional models for diseases like malaria. While the reproduction number in these classical models is very useful for analyzing effects such as reductions due to control measures, it is less flexible for studying virulence evolution. However, similarities can still be drawn between this formulation and more recent ones, such as those proposed by Pinho et al., (2010).
@@
Future studies may explore, within the current framework, the impact of parasite clearance due to immunological effects. The rate of parasite clearance could be integrated into the transmission model as a factor that slows down transmission. This immunological component is expected to become relevant some time after infection, as its effects are minimal during the initial phase. For example, Poehler et al. Poehler et al., (2022) demonstrate the decay of antibodies over time in individuals infected with SARS-CoV-2. Similar antibody decay rates have been observed for DENV (Vaughn et al.,, 2000). While a simple model can yield some insights, subsequent research should consider a more complex model that incorporates immunological factors.
@@
Increased surveillance is essential for vector-borne diseases. The GBD project highlights neglected diseases like leishmaniasis and yellow fever. Pandemic preparedness demands timely data collection and monitoring, especially for currently overlooked diseases. Variability in climate conditions may disrupt existing systems and lead to increased virulence, even if temporary, which can still result in severe cases. The theoretical framework in this study is valuable for advancing knowledge, comparing different scenarios, and assessing risks in changing environments.
@@ Acknowledgments
DAMV is grateful for the National Council for Scientific and Technological Development (CNPq/Brazil) and CAPES (Service Code 001). DAMV is grateful to the Center for Health and Wellbeing (CHW) at Princeton University, as most of this work was done during his time as Visiting Research Scholar at CHW.
@@
The basic reproduction number is maximized when the partial derivatives (2) with respect to vector mortality and host mortality equal zero. Hence, this necessary and sufficient condition permits to find the optimal mortality rates as follows (Obs:maximizing $R_{0}^{2}$).
@@
The left–hand side of Equations 8 and 9 contains the logarithmic derivative (González,, 1992) of functions $\beta_{h}(\omega,\delta)$ and $\beta_{v}(\omega,\delta)$. It is useful to apply the operator $L$, defined by $Lu(x)=\frac{\frac{du}{dx}}{u(x)}$ for function $u(x)$ , and a fraction of $A$ =$f_{A}$. The optimal level of virulence $\omega$ is given by the solution of Equation:
@@
Evaluated diseases in GBD study

### v4

@@ Discussion
The theoretical framework in this study shows that interactions between vector and host transmission position vector-borne diseases within a wide range of virulence states that enable transmission. The vector component may amplify the reproduction number increasing the evolutionary viability for the pathogen and may widen the range of mortality values in which this viability happens. This can lead to increased virulence in hosts, even if it does not correspond necessarily to the highest reproduction numbers. The varying levels of virulence can have serious consequences; for example, yellow fever is known to have a high fatality rate (Douam and Ploss,, 2018). These findings highlight the need for heightened attention to vector-borne diseases as potential emergent threats.
@@
Geographical and climate conditions may restrict vector distribution, meaning they might not be present globally. de Angeli Dutra et al., (2022) also discuss the possibility of higher virulence in vector-host transmission despite the pathogen needing to overcome different immunological systems and the fitness costs to vectors. However, where vectors are present, there is a significant risk of pathogen emergence or adaptation. Therefore, vector-borne diseases pose a high risk, requiring focused surveillance.
@@
The model included a growth rate of pathogens within hosts and vectors in which diminishing transmission rate over virulence levels are crucial. This is biologically plausible since the pathogen will grow in hosts and vectors where there is a finite number of potential cells or tissues to infect. Previous theoretical frameworks have also considered diminishing returns (Dieckmann et al.,, 2002). The tradeoff theory has been discussed by several authors including Ewald, (2004), and a model formulation by Bolker et al., (2009). Recent empirical observations also pointed to diminishing rates for virus titers in SARS-CoV-2 (Mautner et al.,, 2022) in patients and DENV-2 in mosquitoes (Salazar et al.,, 2007). Nevertheless, several authors have also examined the challenges of reproducing such effects in various disease contexts, both experimentally and observationally. Alizon et al., (2009) provide a review discussing why it is often challenging to evaluate the tradeoff hypothesis empirically. In this aspect, Doumayrou et al., (2013) found evidence for the tradeoff hypothesis with plant pathogens, and Jensen et al., (2006) found evidence in experiments with a castrating bacterium. It will be highly valuable if more experimental studies focus on evaluating such conditions for vector-borne diseases.
@@
The initial formulation of the model is quite general to provide intuition on the effects of parameters on the basic reproduction number. However, a few assumptions were applied to obtain some of the theoretical results. Typically, the model assumes that the distributions of general mortality and recovery rates follow an exponential pattern. This assumption simplifies the model, making it easier to analyze and draw meaningful conclusions. The formula for the basic reproduction number in this study is arguably more general than the classical MacDonald and Ross formulation (Smith et al.,, 2012), which is typically used in traditional models for diseases such as malaria and not intended to investigate evolution of virulence. While the derivation of the reproduction number in these classical models is very useful for analyzing effects such as reductions due to control measures, it is less flexible for studying virulence evolution. However, similarities can still be drawn between the present formulation and more recent ones, such as those proposed by Pinho et al., (2010) for the study of dengue virus transmission. Furthermore, the model is sufficiently general to capture transmission cycles of various pathogens, beyond arboviruses and Plasmodium, which use mosquito species as vectors.
@@
Future studies may explore, within the current framework, the impact of parasite clearance due to immunological effects. The rate of parasite clearance could be integrated into the transmission model as a factor that slows down transmission. This immunological component is expected to become relevant sometime after infection, as its effects are minimal during the initial phase. For example, Poehler et al., (2022) demonstrate the decay of antibodies over time in individuals infected with SARS-CoV-2. Similar antibody decay rates have been observed for DENV (Vaughn et al.,, 2000). While a simple model can yield some insights, subsequent research should consider a more complex model that incorporates immunological factors.
@@
Increased surveillance is essential for vector-borne diseases. The GBD project indicates several of these neglected diseases, such as leishmaniasis and yellow fever. Pandemic preparedness demands timely data collection and monitoring, especially for currently overlooked diseases. Variability in climate conditions may disrupt existing systems and lead to increased virulence, even if temporary, which can still result in severe cases. The theoretical framework in this study is valuable for advancing knowledge, comparing different scenarios, and assessing risks in changing environments.
@@ Acknowledgments
DAMV is grateful for the National Council for Scientific and Technological Development (CNPq/Brazil, Ref. 312282/2022-2), Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (Ref. E-26/204.108/2024), and CAPES (Service Code 001). DAMV is grateful to the Center for Health and Wellbeing (CHW) at Princeton University, as most of this work was done during his time as Visiting Research Scholar at CHW.
@@
The basic reproduction number is maximized when the partial derivatives with respect to vector mortality and host mortality equal zero. Hence, this necessary and sufficient condition permits to find the optimal mortality rates as follows (Obs:maximizing $R_{0}^{2}$).
@@
The left–hand side of Equations 8 and 9 contains the logarithmic derivative of functions $\beta_{h}(\omega,\delta)$ and $\beta_{v}(\omega,\delta)$. It is useful to apply the operator $L$, defined by $Lu(x)=\frac{\frac{du}{dx}}{u(x)}$ for function $u(x)$ , and a fraction of $A$ =$f_{A}$. The optimal level of virulence $\omega$ is given by the solution of Equation:
@@
Evaluation of Case Fatality Ratios for countries’ populations
The Global Burden of Disease (GBD) project regularly collects data on various health conditions, including infectious diseases (Institute of Health Metrics and Evaluation,, 2020). The most recent data, as of 2019, provides information on the number of cases and deaths for diseases listed in the Supplementary Text. The study included 169 cause names for diseases and health conditions that might lead to death. A number of 30 infectious diseases were selected from this list with a simple criterion that should involve a pathogen for transmission. Diseases Cisticercosis, nematode infections, Schistosomiasis, and ”other neglected diseases” did not include values for either incidence or mortality and were excluded. The World Bank’s classification of countries per income level was used to categorize countries reported in the study. The final list had 22 infectious diseases or groups of infectious diseases. The case fatality ratio (CFR) of a disease in a given year is calculated in this study by taking the ratio between the incidence and mortality of the disease in that year (Lash et al.,, 2021). This ratio has the advantage of not requiring population estimates for countries.
The three most recent GBD studies were conducted in the years 2019, 2010, and 2000. The fatality rates were observed from empirical data to identify diseases with fatality rates above 1% in recent years. The goal of the original GBD study, however, was not to compare fatality rates among diseases, and some of these diseases may have effective treatment options, such as vaccines.
Figure 5 shows that the diseases from the GBD study (Institute of Health Metrics and Evaluation,, 2020) with higher CFR include vector-host diseases or might have multiple hosts such as schistosomiasis and rabies. As expected, the income level of countries has a noticeable impact, with a higher proportion of diseases exceeding a 1% fatality rate in low-income countries compared to high-income countries. Still, several observations apply to all categories. The variability of case fatality rates of diseases from dengue to leishmaniasis, as listed in Figure 5 is very high. Median values for leishmaniasis, Yellow fever, meningitis, diphteria, tetanus, trypanosomiasis, rabies and schistosomiasis are above 1% for High income countries. The intervals exhibited for Yellow fever, meningitis, diphteria, tetanus, trypanosomiasis, rabies, and schistosomiasis show that above 95% of records are above 1%. Trypanosomiasis, rabies and schostosomiasis have extremely high fatality rates.
List of evaluated diseases in GBD study